acetylene is formed by the reaction of water with calcium carbide according to the following eqation CaC2 +2H2O->Ca(OH)2 +C2H2 .how many grams of CaC2 would be needed to produce 10 L (measured at STP) acetylene .


1
Expert's answer
2022-03-27T10:27:40-0400

CaC2 + 2H2O = Ca(OH)2 + C2H2

n(C2H2) = 10 L / 22.4 L = 0.446 mol

1 mol CaC2 produces 1 mol C2H2

n(CaC2) = 0.446 mol

m(CaC2) = n(CaC2) × Mr(CaC2) = 0.446 mol × 64.099 g/mol = 28.6 g